# Kernel Mean Embedding Topology: Weak and Strong Forms for Stochastic Kernels and Implications for Model Learning

### v1

@@ 3.1 Weak Kernel Mean Embedding Topology on Stochastic Kernels
be a positive semi-definite kernel that induces the reproducing kernel Hilbert space ${\mathcal{H}}_{k}$. We suppose that the following conditions are true for RKHS ${\mathcal{H}}_{k}$ and its kernel $k$. These assumptions are satisfied by well-known kernels like Gaussian, Laplacian, Matern, etc. (see [55, Lemma 8], [14, Proposition 5.6], [59, Section 3.1]).
@@
Under above assumptions, any finite signed measure $\nu\in{\mathcal{M}}({\mathsf{U}})$ can be embedded into ${\mathcal{H}}_{k}$ as follows:
$I:{\mathcal{M}}({\mathsf{U}})\ni\nu\mapsto I_{\nu}\triangleq\int_{{\mathsf{U}}}k(\cdot,u)\,\nu(du)\in{\mathcal{H}}_{k},$
@@
Hence, under this inner product, we can view ${\mathcal{M}}({\mathsf{U}})$ as a pre-Hilbert space whose closure is ${\mathcal{H}}_{k}$. This inner product also induces the following norm, called MMD norm:
$\|\nu_{1}-\nu_{2}\|_{{\mathcal{M}}}\triangleq\sup_{\|h\|_{{\mathcal{H}}_{k}}\leq 1}\left(\int_{{\mathsf{U}}}h(u)\,\nu_{1}(du)-\int_{{\mathsf{U}}}h(u)\,\nu_{2}(du_{2})\right).$
@@
Hence, the MMD norm is weaker than the total variation norm. In fact, on the set of probability measures ${\mathcal{P}}({\mathsf{U}})$, the MMD norm is equivalent to the weak convergence topology, as demonstrated in [55, Lemma 5]. This equivalence implies that the MMD topology is sufficiently weak to establish empirical consistency results as opposed to the total variation norm. Namely, if $\nu$ is a nonatomic measure, then for any singleton $\{u\}$, we have $\nu(\{u\})=0$. This means that for any $n$, if $\nu_{n}$ is the empirical estimate of $\nu$, then the total variation distance $\|\nu_{n}-\nu\|_{TV}$ will always be $2$. Consequently, total variation distance is not suitable for establishing empirical consistency. However, when considering weak convergence topology, the strong law of large numbers ensures that $\nu_{n}\rightarrow\nu$ almost surely as $n\rightarrow\infty$. Therefore, weak convergence topology is the appropriate notion of distance for empirical consistency. Since the MMD topology is equivalent to the weak convergence topology on the set of probability measures, MMD norm as opposed total variation norm can also be effectively used for this purpose. Moreover, the MMD topology endows the set of signed measures with an inner product structure, allowing for the use of geometrical arguments such as orthogonality and projection, which is missing in total variation norm.
@@
$\|\gamma\|_{q}$ =$=\begin{cases}\left(\int_{{\mathsf{Y}}}\|\gamma(y)\|^{q}_{{\mathcal{H}}_{k}}\,\mu(dy)\right)^{1/q},&\,\,\text{if}\,\,q<\infty\\ {\mathrm{ess}}\sup_{y\in{\mathsf{Y}}}\|\gamma(y)\|_{{\mathcal{H}}_{k}},&\,\,\text{if}\,\,q=\infty.\end{cases}$ (4)
@@
that is, we embed $\gamma(y)\in{\mathcal{P}}({\mathsf{U}})$ into ${\mathcal{H}}_{k}$ via $I$ for all $y$.
@@
is also measurable for any $g\in C_{0}({\mathsf{U}})\setminus{\mathcal{H}}_{k}$ as well. Hence, $\gamma$ is a weakly measurable function from ${\mathsf{Y}}$ to ${\mathcal{M}}({\mathsf{U}})$ whose range is contained in ${\mathcal{P}}({\mathsf{U}})$, and so, by Lemma 3.7, $\gamma\in\Gamma$. ∎
Hence, by Lemma 3.1, the set of stochastic kernels can be identified with the following bounded subset $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{P}}})$ of $L_{q}(\mu,{\mathcal{H}}_{k})$:
@@
Let ${\mathcal{H}}_{{\mathcal{P}}_{{}_{\leq 1}}}$ be the image of the set of sub-probability measures ${\mathcal{P}}{{}_{\leq 1}}({\mathsf{U}})$ under $I$. Since ${\mathcal{H}}_{{\mathcal{M}}}$ is a strict subset of ${\mathcal{H}}_{k}$, it is not immediately clear that ${\mathcal{H}}_{{\mathcal{P}}_{{}_{\leq 1}}}$ is a closed subset of ${\mathcal{H}}_{k}$. Similarly, it is not clear that $L_{q}\bigl{(}\mu,{\mathcal{H}}_{{\mathcal{P}}_{{}_{\leq 1}}}\bigr{)}$ is closed in $L_{q}\bigl{(}\mu,{\mathcal{H}}_{k}\bigr{)}$ with respect to the weak kernel mean embedding topology. Therefore, we need to establish these properties first.
@@
First note that if ${\mathsf{U}}$ is compact, then $C_{0}({\mathsf{U}})=C_{b}({\mathsf{U}})$. Let $\gamma_{\lambda}\rightharpoonup^{*}\gamma$ with respect to weak kernel mean embedding topology. Since $L_{q}\bigl{(}\mu,{\mathcal{H}}_{{\mathcal{P}}_{{}_{\leq 1}}}\bigr{)}$ is closed in $L_{q}\bigl{(}\mu,{\mathcal{H}}_{k}\bigr{)}$ with respect to weak kernel mean embedding topology, we have $\gamma\in L_{q}\bigl{(}\mu,{\mathcal{H}}_{{\mathcal{P}}_{{}_{\leq 1}}}\bigr{)}$. We now prove that $\gamma$ is indeed an element of $L_{q}\bigl{(}\mu,{\mathcal{H}}_{{\mathcal{P}}}\bigr{)}$. By $\gamma_{\lambda}\rightharpoonup^{*}\gamma$, we have $\mu\otimes\gamma_{\lambda}({\mathsf{Y}}\times\cdot)\rightarrow\mu\otimes\gamma({\mathsf{Y}}\times\cdot)$ in weak MMD topology [^3] A sequence $\{\nu_{n}\}\in{\mathcal{M}}({\mathsf{U}})$ converges to $\nu$ in weak MMD topology if $\langle{\mathcal{I}}_{\nu_{n}},h\rangle_{{\mathcal{H}}_{k}}=\int_{{\mathsf{U}}}h(u)\,\nu_{n}(du)\rightarrow\langle{\mathcal{I}}_{\nu},h\rangle_{{\mathcal{H}}_{k}}=\int_{{\mathsf{U}}}h(u)\,\nu(du)\,\,\forall h\in{\mathcal{H}}_{k}.$ The relationship between the MMD topology and the weak MMD topology is very similar to the relationship between the topology induced by the total variation norm and the weak convergence topology. In the former case, the convergence of the integral is uniform over the unit ball in the function class of interest, while in the latter case, the convergence is pointwise within that function class.; that is,
@@
By [55, Lemma 5 and Remark 6], $\mu\otimes\gamma_{\lambda}({\mathsf{Y}}\times\cdot)\rightarrow\mu\otimes\gamma({\mathsf{Y}}\times\cdot)$ in vague topology, and so, in weak convergence topology as $C_{0}({\mathsf{U}})=C_{b}({\mathsf{U}})$. Since $\mu\otimes\gamma_{\lambda}({\mathsf{Y}}\times\cdot)\in{\mathcal{P}}({\mathsf{U}})$ for all $\lambda$ and ${\mathcal{P}}({\mathsf{U}})$ is a closed subset of ${\mathcal{M}}({\mathsf{U}})$ under weak convergence topology, we have $\mu\otimes\gamma({\mathsf{Y}}\times\cdot)\in{\mathcal{P}}({\mathsf{U}})$. Hence, $\gamma\in\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{P}}({\mathsf{U}})\bigr{)}$.
Suppose now that ${\mathsf{U}}$ is not compact. Using the same construction as in the proof of Lemma 3.8, we can find a sequence of probability measures $\{\nu_{n}\}$ such that $\nu_{n}$ converges to 00 vaguely, where $0(\,\cdot\,)$ denotes the degenerate measure on ${\mathsf{U}}$. Then, define $\gamma_{n}(y)(\,\cdot\,)=\nu_{n}(\,\cdot\,)$ and $\gamma(y)(\,\cdot\,)=0(\,\cdot\,)$. Let $f\in L_{p}\bigl{(}\mu,{\mathcal{H}}_{k}\bigr{)}$. Then we have
@@
[…]
@@
In view of Lemma 3.4, $L_{q}\bigl{(}\mu,{\mathcal{H}}_{{\mathcal{P}}}\bigr{)}$ is relatively (sequentially) compact with respect to weak kernel mean embedding topology.
@@
The following result establishes that convergence under the weak kernel mean embedding topology remains valid even if the reference probability measure $\mu$ on ${\mathsf{Y}}$ is replaced by another probability measure $\eta$, provided $\eta$ is absolutely continuous with respect to $\mu$. A similar result is established in [68, Lemma 3.6] for Young narrow topology.
@@
Let $\eta\ll\mu$, and suppose the Radon-Nikodym derivative of $\eta$ with respect to $\mu$ is $l$-integrable with respect to $\mu$. Then, $\gamma_{n}\rightharpoonup^{*}\gamma$ under the weak kernel mean embedding topology at input $\mu$ implies $\gamma_{n}\rightharpoonup^{*}\gamma$ under the weak kernel mean embedding topology at input $\eta$, provided $p$ is replaced by $\tilde{p}$ satisfying $1/\tilde{p}+1/l=1/p$. Furthermore, if $l>1$, since the weak kernel mean embedding topologies are equivalent for any $1\leq p<\infty$, there is no need to adjust $p$ in this case.
@@
$\bigg{|}\int_{{\mathsf{Y}}}\|\tilde{g}(y)\|_{{\mathcal{H}}_{k}}^{p}\mu(dy)\bigg{|}^{1/p}\leq\|f\|_{l}\,\|g\|_{\tilde{p}}<\infty.$

### v2

@@ 3.1 Weak Kernel Mean Embedding Topology on Stochastic Kernels
be a positive semi-definite kernel that induces the reproducing kernel Hilbert space ${\mathcal{H}}_{k}$. Indeed, let us define the feature function $\Phi$ as
$\Phi(u)\triangleq k(\cdot,u),\,\,\forall u\in{\mathsf{U}}.$
Therefore, we have
${\mathcal{H}}_{k}=\mathop{\rm cl}\mathop{\rm span}\{\Phi(u):u\in{\mathsf{U}}\},$
where the closure is taken with respect to the topology induced by the inner product
$\left\langle\sum_{i=1}^{n}\alpha_{i}\,\Phi(u_{i}),\sum_{j=1}^{m}\gamma_{j}\,\Phi(u_{j})\right\rangle_{{\mathcal{H}}_{k}}\triangleq\sum_{i=1,j=1}^{n,m}\alpha_{i}\,\gamma_{j}\,k(u_{i},u_{j}).$
This inner product has the reproducing property, that is, for any $f\in{\mathcal{H}}_{k}$, we have
$f(u)=\langle f,\Phi(u)\rangle_{{\mathcal{H}}_{k}}.$
In particular, this implies
$k(u_{1},u_{2})=\langle\Phi(u_{1}),\Phi(u_{2})\rangle_{{\mathcal{H}}_{k}},$
which suggests that the kernel $k$ can be interpreted as an inner product between two feature mappings in $\mathcal{H}_{k}$.
We suppose that the following conditions are true for RKHS ${\mathcal{H}}_{k}$ and its kernel $k$. These assumptions are satisfied by well-known kernels like Gaussian, Laplacian, Matern, etc. (see [67, Lemma 8], [22, Proposition 5.6], [71, Section 3.1]).
@@
Under the assumptions above, any finite signed measure $\nu\in{\mathcal{M}}({\mathsf{U}})$ can be embedded into ${\mathcal{H}}_{k}$ as follows:
$I:{\mathcal{M}}({\mathsf{U}})\ni\nu\mapsto I_{\nu}\triangleq\int_{{\mathsf{U}}}k(\cdot,u)\,\nu(du)\in{\mathcal{H}}_{k},$ (3)
@@
where we use the fundamental RKHS property that $\langle k(\cdot,u_{1}),k(\cdot,u_{2})\rangle_{{\mathcal{H}}_{k}}=k(u_{1},u_{2})$. Hence, under this inner product, we can view ${\mathcal{M}}({\mathsf{U}})$ as a pre-Hilbert space whose closure is ${\mathcal{H}}_{k}$. This inner product also induces the following norm, called MMD norm:
$\|\nu_{1}-\nu_{2}\|_{{\mathcal{M}}}$ $\triangleq\|I_{\nu_{1}}-I_{\nu_{2}}\|_{{\mathcal{H}}_{k}}$
$=\sup_{\|h\|_{{\mathcal{H}}_{k}}\leq 1}\langle h,I_{\nu_{1}}-I_{\nu_{2}}\rangle_{{\mathcal{H}}_{k}}$
$=\sup_{\|h\|_{{\mathcal{H}}_{k}}\leq 1}\left(\langle h,I_{\nu_{1}}\rangle_{{\mathcal{H}}_{k}}-\langle h,I_{\nu_{2}}\rangle_{{\mathcal{H}}_{k}}\right)$
$=\sup_{\|h\|_{{\mathcal{H}}_{k}}\leq 1}\left(\left\langle h,\int_{{\mathsf{U}}}k(\cdot,u)\,\nu_{1}(du)\right\rangle_{{\mathcal{H}}_{k}}-\left\langle h,\int_{{\mathsf{U}}}k(\cdot,u)\,\nu_{2}(du)\right\rangle_{{\mathcal{H}}_{k}}\right)$
$=\sup_{\|h\|_{{\mathcal{H}}_{k}}\leq 1}\left(\int_{{\mathsf{U}}}\langle h,k(\cdot,u)\rangle_{{\mathcal{H}}_{k}}\,\nu_{1}(du)-\int_{{\mathsf{U}}}\langle h,k(\cdot,u)\rangle_{{\mathcal{H}}_{k}}\,\nu_{2}(du)\right)$
$=\sup_{\|h\|_{{\mathcal{H}}_{k}}\leq 1}\left(\int_{{\mathsf{U}}}h(u)\,\nu_{1}(du)-\int_{{\mathsf{U}}}h(u)\,\nu_{2}(du_{2})\right).$
@@
Hence, the MMD norm is weaker than the total variation norm. In fact, on the set of probability measures ${\mathcal{P}}({\mathsf{U}})$, the MMD norm is equivalent to the weak convergence topology, as demonstrated in [67, Lemma 5]. This equivalence implies that the MMD topology is sufficiently weak to establish empirical consistency results as opposed to the total variation norm. Namely, if $\nu$ is a nonatomic measure, then for any singleton $\{u\}$, we have $\nu(\{u\})=0$. This means that for any $n$, if $\nu_{n}$ is the empirical estimate of $\nu$, then the total variation distance $\|\nu_{n}-\nu\|_{TV}$ will always be $2$. Consequently, total variation distance is not suitable for establishing empirical consistency. However, when considering weak convergence topology, the strong law of large numbers ensures that $\nu_{n}\rightarrow\nu$ almost surely as $n\rightarrow\infty$. Therefore, weak convergence topology is the appropriate notion of distance for empirical consistency. Since the MMD topology is equivalent to the weak convergence topology on the set of probability measures, MMD norm as opposed total variation norm can also be effectively used for this purpose. Moreover, the MMD topology endows the set of signed measures with an inner product structure, allowing for the use of geometrical arguments such as orthogonality and projection, which is missing in total variation norm.
@@
$\|\gamma\|_{q}$ $=\begin{cases}\left(\int_{{\mathsf{Y}}}\|\gamma(y)\|^{q}_{{\mathcal{H}}_{k}}\,\mu(dy)\right)^{1/q},&\,\,\text{if}\,\,q<\infty\\ {\mathrm{ess}}\sup_{y\in{\mathsf{Y}}}\|\gamma(y)\|_{{\mathcal{H}}_{k}},&\,\,\text{if}\,\,q=\infty.\end{cases}$ (5)
@@
that is, we embed $\gamma(y)\in{\mathcal{P}}({\mathsf{U}})$ into ${\mathcal{H}}_{k}$ via $I$ for all $y$ (recall the definition of $I$ in (3)).
@@
is also measurable for any $g\in C_{0}({\mathsf{U}})\setminus{\mathcal{H}}_{k}$ as well. Hence, $\gamma$ is a weakly measurable function from $\mathsf{Y}$} to ${\mathcal{M}}({\mathsf{U}})$ whose range is contained in ${\mathcal{P}}({\mathsf{U}})$. As Lemma 3.7 will show, $\Gamma$ consists exactly of those weakly measurable functions from $\mathsf{Y}$} to ${\mathcal{M}}({\mathsf{U}})$ whose range lies in ${\mathcal{P}}({\mathsf{U}})$.
Hence, by Lemma 3.1, the set of stochastic kernels can be identified with the following bounded subset $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{P}}})$ of $L_{q}(\mu,{\mathcal{H}}_{k})$ (Here, we let $q\in(1,\infty]$, though in some parts we exclude the case $q=\infty$ in order to make use of reflexivity. The case $q=2$ is of particular interest because of the underlying Hilbert space structure.):
@@
[…]
@@
First note that if $\mathsf{U}$} is compact, then $C_{0}({\mathsf{U}})=C_{b}({\mathsf{U}})=C({\mathsf{U}})$. Let $\gamma_{\lambda}\rightharpoonup^{*}\gamma$ with respect to weak kernel mean embedding topology. Since $L_{q}\bigl(\mu,{\mathcal{H}}_{{\mathcal{P}}_{{}_{\leq 1}}}\bigr)$ is closed in $L_{q}\bigl(\mu,{\mathcal{H}}_{k}\bigr)$ with respect to weak kernel mean embedding topology, we have $\gamma\in L_{q}\bigl(\mu,{\mathcal{H}}_{{\mathcal{P}}_{{}_{\leq 1}}}\bigr)$. We now prove that $\gamma$ is indeed an element of $L_{q}\bigl(\mu,{\mathcal{H}}_{{\mathcal{P}}}\bigr)$. By $\gamma_{\lambda}\rightharpoonup^{*}\gamma$, we have $\mu\otimes\gamma_{\lambda}({\mathsf{Y}}\times\cdot)\rightarrow\mu\otimes\gamma({\mathsf{Y}}\times\cdot)$ in weak MMD topology [^3] A sequence $\{\nu_{n}\}\in{\mathcal{M}}({\mathsf{U}})$ converges to $\nu$ in weak MMD topology if $\langle{\mathcal{I}}_{\nu_{n}},h\rangle_{{\mathcal{H}}_{k}}=\int_{{\mathsf{U}}}h(u)\,\nu_{n}(du)\rightarrow\langle{\mathcal{I}}_{\nu},h\rangle_{{\mathcal{H}}_{k}}=\int_{{\mathsf{U}}}h(u)\,\nu(du)\,\,\forall h\in{\mathcal{H}}_{k}.$ The relationship between the MMD topology and the weak MMD topology is very similar to the relationship between the topology induced by the total variation norm and the weak convergence topology. In the former case, the convergence of the integral is uniform over the unit ball in the function class of interest, while in the latter case, the convergence is pointwise within that function class.; that is,
@@
By [67, Lemma 5 and Remark 6], $\mu\otimes\gamma_{\lambda}({\mathsf{Y}}\times\cdot)\rightarrow\mu\otimes\gamma({\mathsf{Y}}\times\cdot)$ in vague topology, and so, in weak convergence topology as $C_{0}({\mathsf{U}})=C_{b}({\mathsf{U}})=C({\mathsf{U}})$. Since $\mu\otimes\gamma_{\lambda}({\mathsf{Y}}\times\cdot)\in{\mathcal{P}}({\mathsf{U}})$ for all $\lambda$ and ${\mathcal{P}}({\mathsf{U}})$ is a closed subset of ${\mathcal{M}}({\mathsf{U}})$ under weak convergence topology, we have $\mu\otimes\gamma({\mathsf{Y}}\times\cdot)\in{\mathcal{P}}({\mathsf{U}})$. Hence, $\gamma\in\mathcal{L}_{q}\bigl(\mu,{\mathcal{P}}({\mathsf{U}})\bigr)$.
Suppose now that $\mathsf{U}$} is not compact. Using the same construction as in the proof of Lemma 3.8, we can find a sequence of probability measures $\{\nu_{n}\}$ such that $\nu_{n}$ converges to 0 vaguely, where $0(\,\cdot\,)$ denotes the degenerate measure on $\mathsf{U}$}. Then, define $\gamma_{n}(y)(\,\cdot\,)=\nu_{n}(\,\cdot\,)$ and $\gamma(y)(\,\cdot\,)=0(\,\cdot\,)$. Let $f\in L_{p}\bigl(\mu,{\mathcal{H}}_{k}\bigr)$. Then we have
@@
$=\int_{{\mathsf{Y}}}\lim_{n\rightarrow\infty}\int_{{\mathsf{U}}}f(y)(u)\,\nu_{n}(du)\,\mu(dy)\text{ (as $\|f(y)\|_{{\mathcal{H}}_{k}}$ is $\mu$-integrable)}$
$=0\text{ (as $f(y)\in{\mathcal{H}}_{k}\subset C_{0}({\mathsf{U}})$)}.$
@@
Recall that $L_{q}\bigl(\mu,{\mathcal{H}}_{\mathcal{P}_{\leq 1}}\bigr)$ is (sequentially) compact with respect to the weak kernel mean embedding topology by the Banach–Alaoglu Theorem. Since $L_{q}\bigl(\mu,{\mathcal{H}}_{\mathcal{P}}\bigr)$ is a subset of $L_{q}\bigl(\mu,{\mathcal{H}}_{\mathcal{P}_{\leq 1}}\bigr)$, it is relatively (sequentially) compact under the same topology. Moreover, when $\mathsf{U}$} is compact, this set is also (sequentially) compact by Lemma 3.4. In contrast, when $\mathsf{U}$} is non-compact, Lemma 3.4 shows that $L_{q}\bigl(\mu,{\mathcal{H}}_{\mathcal{P}}\bigr)$ is no longer closed, and therefore compactness is not guaranteed.
@@
The following result establishes that convergence under the weak kernel mean embedding topology remains valid even if the reference probability measure $\mu$ on $\mathsf{Y}$} is replaced by another probability measure $\eta$, provided $\eta$ is absolutely continuous with respect to $\mu$. A similar result is established in [79, Lemma 3.6] for Young narrow topology.
@@
Let $\eta\ll\mu$, and suppose the Radon-Nikodym derivative of $\eta$ with respect to $\mu$ is $l$-integrable with respect to $\mu$, where $l>1$. Then, $\gamma_{n}\rightharpoonup^{*}\gamma$ under the weak kernel mean embedding topology at input $\mu$ implies $\gamma_{n}\rightharpoonup^{*}\gamma$ under the weak kernel mean embedding topology at input $\eta$, provided $p$ is replaced by $\tilde{p}$ satisfying $1/\tilde{p}+1/l=1/p$. Indeed, since the weak kernel mean embedding topologies are equivalent for any $1\leq p<\infty$, there is no need to adjust $p$.
@@
$\bigg|\int_{{\mathsf{Y}}}\|\tilde{g}(y)\|_{{\mathcal{H}}_{k}}^{p}\mu(dy)\bigg|^{1/p}\leq\|f\|_{l}\,\|g\|_{\tilde{p}}$$\infty$.